OpenStudy (blackops2_143):

I'm having some trouble understanding this question. please help Solve for z. z−49−13=59

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (blackops2_143):

it's supposed to look like this oops z−4/9−1/3=5/9

OpenStudy (skullpatrol):

Ok, what do you think?

OpenStudy (blackops2_143):

do I need to add to both sides?

OpenStudy (skullpatrol):

Yup...

OpenStudy (blackops2_143):

okay, do I add 1/3 to both sides?

OpenStudy (skullpatrol):

You could, but look what you're adding to...

OpenStudy (skullpatrol):

It has a 9 in the denominator, right?

OpenStudy (blackops2_143):

yeah

OpenStudy (blackops2_143):

it's larger than the denominator?

OpenStudy (skullpatrol):

what do you mean?

OpenStudy (blackops2_143):

oops numerator

OpenStudy (blackops2_143):

the denominator is larger than the numerator so it won't work right?

OpenStudy (skullpatrol):

It will work if you make them have the SAME denominators

OpenStudy (blackops2_143):

so add 4 to the 5

OpenStudy (skullpatrol):

yup

OpenStudy (blackops2_143):

ok

OpenStudy (blackops2_143):

so z - 9/9 - 1/3?

OpenStudy (skullpatrol):

nope

OpenStudy (blackops2_143):

um okay. any suggestions then?

OpenStudy (blackops2_143):

I'm not fully understanding

OpenStudy (skullpatrol):

let's start again

OpenStudy (skullpatrol):

z−4/9−1/3=5/9

OpenStudy (blackops2_143):

k so should I add 4/9 to 5/9? or what should I do there?

OpenStudy (skullpatrol):

yes z−4/9−1/3=5/9 4/9 + z − 4/9 − 1/3 = 5/9 + 4/9

OpenStudy (skullpatrol):

add 4/9 to BOTH sides

OpenStudy (blackops2_143):

ok thank you

OpenStudy (skullpatrol):

what do you get after adding 4/9 to both sides?

OpenStudy (blackops2_143):

z = 4/3

OpenStudy (skullpatrol):

Good, now take that value of z and substitute it into the original equation: z−4/9−1/3=5/9 to check your work

OpenStudy (blackops2_143):

it's right

OpenStudy (skullpatrol):

Good job :D

OpenStudy (blackops2_143):

thank you!!